# Sobolev regularity theory for stochastic reaction-diffusion-advection equations with spatially homogeneous colored noises and variable-order nonlocal operators

### v2

@@ 2.4. Solvability of semilinear stochastic integro-diffrential equations
$-\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left(\overline{v(x)}-\overline{v(y)}\right)\left(u(x)-u(y)\right)j(|x-y|)\mathrm{d}x\mathrm{d}y.$
@@
where $N=N()$. For achieving (3.17), the remaining part of step 1 is only to prove
@@
converges to 00 as $\varepsilon_{1},\varepsilon_{2}\downarrow 0$. Due to (3.18), it can be easily checked that $\{u_{1/m}\}_{m=1}^{\infty}$ is a Cauchy sequence in $L_{2}(\Omega;L_{\infty}([0,\tau];L_{2}(\mathbb{R}^{d})))$. Therefore, there exists a limit $\bar{u}$ such that $u_{1/m}\to\bar{u}$ in $L_{2}(\Omega;L_{\infty}([0,\tau];L_{2}(\mathbb{R}^{d})))$. One can also easily check that $u_{1/m}(t,\cdot)\to u(t,\cdot)$ in the sense of distribution for all $t\leq\tau$ almost surely. Therefore, with probability $1$, $u(t,\cdot)\in L_{2}(\mathbb{R}^{d})$ for all $t\leq\tau$. This certainly implies that $\bar{u}(t,\cdot)=u(t,\cdot)$ in $L_{2}(\mathbb{R}^{d})$ for all $t\leq\tau$, thus (3.19) holds.
@@
$2\int_{\mathbb{R}^{d}}u_{+}(s,x)\phi(\Delta)u(s,x)\mathrm{d}x$ (3.21)
$=-\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}(u_{+}(s,x)-u_{+}(s,y))(u(s,x)-u(s,y))\mathrm{d}x\mathrm{d}y\leq 0.$
Merging (3.20) and (3.21), (3.16) is thus affirmed. Thus, the theorem is conclusively proved. ∎
@@
$f_{m}^{n}:=(\zeta F_{m}(u_{m}^{n})+\vec{b}\cdot\nabla_{x}(B_{m}(u_{m}^{n})))1_{\text{$[$\kern-2.66003pt$[$}0,\tau\text{$]$\kern-1.04996pt$]$}}\in L_{2}(\Omega\times(0,\infty)\times\mathbb{R}^{d}),$
@@
$g_{m}^{n}:=\xi\varphi_{m}(u_{m}^{n})\bm{\pi}_{n}1_{\text{$[$\kern-2.66003pt$[$}0,\tau\text{$]$\kern-1.04996pt$]$}}\in L_{2}(\Omega\times(0,\infty)\times\mathbb{R}^{d};\ell_{2}),\quad\bm{\pi}_{n}:=(\pi\ast e_{1},\cdots,\pi\ast e_{n},0,\cdots),$
@@
The proof of Theorem 3.1-(iii), which addresses the space-time Hölder type regularity of local solutions, follows directly as a corollary from the theorem below.
@@
Given $p\in[2,\infty)$, $q\in(2,\infty)$, and $\gamma\in(0,\infty)$, under Assumption 1.11, for
$\frac{1}{q}\leq\alpha<\beta\leq\frac{1}{2},\quad\delta_{0}(\gamma-2\beta)-\frac{d}{p}>0,$
the space $\mathcal{H}_{p,q}^{\phi,\gamma}(T)$ is continuously embedded into $L_{q}(\Omega;C_{t,x}^{\alpha-\frac{1}{q},\phi_{\gamma-2\beta}-\frac{d}{p}}([0,T]\times\mathbb{R}^{d}))$. In other words, there exists a constant $N=N(d,N_{0},N_{1},N_{2},p,q,T)$ such that
@@
Here, the Hölder norm $|u|_{C_{t,x}^{\alpha-\frac{1}{q},\phi_{\gamma-2\beta}-\frac{d}{p}}([0,T]\times\mathbb{R}^{d})}$ is defined as
$|u|_{C_{t,x}^{\alpha-\frac{1}{q},\phi_{\gamma-2\beta}-\frac{d}{p}}([0,T]\times\mathbb{R}^{d})}:=|u|_{C([0,T]\times\mathbb{R}^{d})}+\sup_{|h_{1}|\leq T,|h_{2}|\leq 1}\frac{|\mathcal{D}_{(h_{1},h_{2})}^{L}v|_{C([0,T]\times\mathbb{R}^{d})}}{|h_{1}|^{\alpha-\frac{1}{q}}+\phi(|h_{2}|^{-2})^{-\gamma+2\beta}|h_{2}|^{-\frac{d}{p}}},$
$\mathcal{D}_{(h_{1},h_{2})}^{L}v(t,x):=\mathcal{D}_{(h_{1},h_{2})}(\mathcal{D}_{(h_{1},h_{2})}^{L-1}v)(t,x)$, $\mathcal{D}_{(h_{1},h_{2})}v(t,x):=v(t+h_{1},x+h_{2})-v(t,x)$ and $L$ is a natural number greater than $\gamma-d/p$.

### v3

@@ 2.4. Solvability of semilinear stochastic integro-differential equations
$\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left(\overline{v(x)}-\overline{v(y)}\right)\left(u(x)-u(y)\right)j(|x-y|)\mathrm{d}x\mathrm{d}y-b_{\phi}\int_{\mathbb{R}^{d}}\overline{\nabla v(x)}\cdot\nabla u(x)\mathrm{d}x.$
@@
where $N=N(d,\delta,\lambda,T)$. For achieving (3.16), the remaining part of step 1 is only to prove
@@
converges to 0 as $\varepsilon_{1},\varepsilon_{2}\downarrow 0$. Due to (3.17), it can be easily checked that $\{u_{1/m}\}_{m=1}^{\infty}$ is a Cauchy sequence in $L_{2}(\Omega;L_{\infty}([0,\tau];L_{2}(\mathbb{R}^{d})))$. Therefore, there exists a limit $\bar{u}$ such that $u_{1/m}\to\bar{u}$ in $L_{2}(\Omega;L_{\infty}([0,\tau];L_{2}(\mathbb{R}^{d})))$. One can also easily check that $u_{1/m}(t,\cdot)\to u(t,\cdot)$ in the sense of distribution for all $t\leq\tau$ almost surely. Therefore, with probability $1$, $u(t,\cdot)\in L_{2}(\mathbb{R}^{d})$ for all $t\leq\tau$. This certainly implies that $\bar{u}(t,\cdot)=u(t,\cdot)$ in $L_{2}(\mathbb{R}^{d})$ for all $t\leq\tau$, thus (3.18) holds.
@@
$2\int_{\mathbb{R}^{d}}u_{+}(s,x)\phi(\Delta)u(s,x)\mathrm{d}x$
$=-\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}(u_{+}(s,x)-u_{+}(s,y))(u(s,x)-u(s,y))j(|x-y|)\mathrm{d}x\mathrm{d}y$
$\quad-2b_{\phi}\int_{\mathbb{R}^{d}}\nabla u_{+}(x)\nabla u(x)\mathrm{d}x.$
Observe that
$\big{(}u_{+}(s,x)-u_{+}(s,y)\big{)}\big{(}u(s,x)-u(s,y)\big{)}\geq 0,$
and for almost every $x\in\mathbb{R}^{d}$,
$\nabla u_{+}(x)=1_{\{u>0\}}(x)\nabla u(x).$
It follows that
$2\int_{\mathbb{R}^{d}}u_{+}(s,x)\,\phi(\Delta)u(s,x)\,\mathrm{d}x\leq 0.$ (3.20)
Combining (3.19) and (3.20), we obtain the desired inequality (3.15). This completes the proof of the theorem. ∎
Remark 3.6.
Theorem 3.5 is stated as a nonpositivity principle. The nonnegativity version is equivalent by a sign change. Indeed, let $u$ solve
$\mathrm{d}u=(\phi(\Delta)u+f)\,\mathrm{d}t+\sum_{k=1}^{\infty}g^{k}\,\mathrm{d}w_{t}^{k},$
and assume
$u(0,\cdot)\geq 0,\qquad 1_{\{u<0\}}\,f\geq 0,\qquad 1_{\{u<0\}}\,g^{k}=0.$
Set v:=uv:=-u. Then $v$ solves
$\mathrm{d}v=(\phi(\Delta)v-f)\,\mathrm{d}t-\sum_{k=1}^{\infty}g^{k}\,\mathrm{d}w_{t}^{k},$
and satisfies
$v(0,\cdot)\leq 0,\qquad 1_{\{v>0\}}(-f)=-1_{\{u<0\}}f\leq 0,\qquad 1_{\{v>0\}}(-g^{k})=-1_{\{u<0\}}g^{k}=0.$
Applying Theorem 3.5 to $v$ yields $v\leq 0$, hence $u\geq 0$. Thus, the nonpositivity and nonnegativity formulations are equivalent.
Remark 3.7.
Theorem 3.5 says the nonpositivity of the solution $u$. However, the nonpositivity principle and the nonnegativity principle are equivalent. Suppose that $u$ is a solution to
Suppose also that
$u(0,\cdot)\geq 0,\quad 1_{u(t,\cdot)<0}f(t,\cdot)\geq 0,\quad 1_{u(t,\cdot)<0}g^{k}(t,\cdot)=0.$
Then v:=uv:=-u is a solution to
$\mathrm{d}v=(\phi(\Delta)v-f)\mathrm{d}t-\sum_{k=1}^{\infty}g^{k}\mathrm{d}w^{k}_{t}.$
Then
$v(0,\cdot)=-u(0,\cdot)\leq 0,\quad 1_{v(t,\cdot)>0}(-f(t,\cdot))=-1_{u(t,\cdot)<0}f(t,\cdot)\leq 0,$
and
$1_{v(t,\cdot)>0}(-g^{k}(t,\cdot))=-1_{u(t,\cdot)<0}g^{k}(t,\cdot)=0.$
By Theorem 3.5, we have $u=v\leq 0$, thus $u\geq 0$.
@@
$f_{m}^{n}:=(\zeta F_{m}(u_{m}^{n})+\vec{b}\cdot\nabla_{x}(B_{m}(u_{m}^{n})))1_{\text{$[$\kern-3.02948pt$[$}0,\tau\text{$]$\kern-1.19579pt$]$}}\in L_{2}(\Omega\times(0,\infty)\times\mathbb{R}^{d}),$
@@
$g_{m}^{n}:=\xi\varphi_{m}(u_{m}^{n})\bm{\pi}_{n}1_{\text{$[$\kern-3.02948pt$[$}0,\tau\text{$]$\kern-1.19579pt$]$}}\in L_{2}(\Omega\times(0,\infty)\times\mathbb{R}^{d};\ell_{2}),\quad\bm{\pi}_{n}:=(\pi\ast e_{1},\cdots,\pi\ast e_{n},0,\cdots),$
@@
The proof of Theorem 3.1-(iii), which concerns the space–time Hölder-type regularity of local solutions, follows as an immediate corollary of the theorem stated below. We begin by recalling the definition of parabolic Hölder spaces adapted to our setting.
Definition 3.8 (Parabolic Hölder space).
Let $\delta_{0}\in(0,1]$, $\gamma\in(0,\infty)$, $p\in[2,\infty)$, $q\in(2,\infty)$, and let $\phi$ be a Bernstein function satisfying Assumption 1.12 ($\delta_{0}$). Assume that
$\frac{1}{q}\leq\alpha<\beta\leq\frac{1}{2},\qquad\delta_{0}(\gamma-2\beta)-\frac{d}{p}>0.$
For an integer $L>\gamma-\frac{d}{p}$, define the $L$-th order difference in the parabolic direction $(\tau,h)\in\mathbb{R}\times\mathbb{R}^{d}$ by
$\mathcal{D}^{0}_{(\tau,h)}u:=u,\qquad\mathcal{D}^{L}_{(\tau,h)}u(t,x):=\mathcal{D}^{1}_{(\tau,h)}\mathcal{D}^{L-1}_{(\tau,h)}u(t,x)=\sum_{k=0}^{L}(-1)^{L-k}\binom{L}{k}\,u(t+k\tau,x+kh).$
For $T>0$, set
$\mathbb{R}^{d}_{T,L,\tau}:=\Big{\{}(t,x)\in[0,T]\times\mathbb{R}^{d}:\ (t+L\tau,x)\in[0,T]\times\mathbb{R}^{d}\Big{\}}.$
We denote by $C_{t,x}^{\alpha-\frac{1}{q},\,\phi_{\gamma-2\beta}-\frac{d}{p}}([0,T]\times\mathbb{R}^{d})$ the set of continuous functions $u$ on $0,T]\times\mathbb{R}^{d}$ such that the norm
$|u|_{C_{t,x}^{\alpha-\frac{1}{q},\,\phi_{\gamma-2\beta}-\frac{d}{p}}([0,T]\times\mathbb{R}^{d})}:=|u|_{C([0,T]\times\mathbb{R}^{d})}+\sup_{(\tau,h)\neq(0,0)}\frac{\|\mathcal{D}^{L}_{(\tau,h)}u\|_{L_{\infty}(\mathbb{R}^{d}_{T,L,\tau})}}{|\tau|^{\alpha-\frac{1}{q}}+\phi(|h|^{-2})^{-\frac{\gamma}{2}+\beta}\,|h|^{-\frac{d}{p}}}$
is finite.
Proposition 3.9.
The space $C_{t,x}^{\alpha-\frac{1}{q},\,\phi_{\gamma-2\beta}-\frac{d}{p}}([0,T]\times\mathbb{R}^{d})$ can be characterized as follows:
$C_{t,x}^{\alpha-\frac{1}{q},\,\phi_{\gamma-2\beta}-\frac{d}{p}}([0,T]\times\mathbb{R}^{d})=C(\mathbb{R}^{d};C^{\alpha-\frac{1}{q}}([0,T]))\cap C([0,T];C^{\phi_{\gamma-2\beta}-\frac{d}{p}}(\mathbb{R}^{d})),$
where
$|u|_{C(\mathbb{R}^{d};C^{\alpha-\frac{1}{q}}([0,T]))}=|u|_{C([0,T]\times\mathbb{R}^{d})}+\sup_{x\in\mathbb{R}^{d}}\sup_{t,s\in[0,T]}\frac{|u(t,x)-u(s,x)|}{|t-s|^{\alpha-\frac{1}{q}}},$
$|u|_{C([0,T];C^{\phi_{\gamma-2\beta}-\frac{d}{p}}(\mathbb{R}^{d}))}=|u|_{C([0,T]\times\mathbb{R}^{d})}+\sup_{t\in[0,T]}\sup_{\begin{subarray}{c}x\in\mathbb{R}^{d}\\ |h|\leq 1\end{subarray}}\frac{|\mathcal{D}^{L}_{(0,h)}u(t,x)|}{\phi(|h|^{-2})^{-\frac{\gamma}{2}+\beta}\,|h|^{-\frac{d}{p}}}.$
Proof.
First, note that
$|\mathcal{D}^{L}_{(\tau,h)}u(t,x)|$ $\leq|\mathcal{D}^{L}_{(\tau,h)}u(t,x)-\mathcal{D}^{L}_{(0,h)}u(t,x)|+|\mathcal{D}^{L}_{(0,h)}u(t,x)|$
$\leq N(L)\sum_{j=1}^{L}|u(t+j\tau,x+jh)-u(t,x+jh)|+|\mathcal{D}^{L}_{(0,h)}u(t,x)|.$
Hence
$\frac{\|\mathcal{D}^{L}_{(\tau,h)}u\|_{L_{\infty}(\mathbb{R}^{d}_{T,L,\tau})}}{|\tau|^{\alpha-\frac{1}{q}}+\phi(|h|^{-2})^{-\frac{\gamma}{2}+\beta}|h|^{-\frac{d}{p}}}$
$\quad\leq N(\alpha,L,q)\sup_{x\in\mathbb{R}^{d}}\sup_{t,s\in[0,T]}\frac{|u(t,x)-u(s,x)|}{|t-s|^{\alpha-\frac{1}{q}}}+\sup_{t\in[0,T]}\sup_{\begin{subarray}{c}x\in\mathbb{R}^{d}\\ |h|\leq 1\end{subarray}}\frac{|\mathcal{D}^{L}_{(0,h)}u(t,x)|}{\phi(|h|^{-2})^{-\frac{\gamma}{2}+\beta}|h|^{-\frac{d}{p}}}.$
This shows that
$C(\mathbb{R}^{d};C^{\alpha-\frac{1}{q}}([0,T]))\cap C([0,T];C^{\phi_{\gamma-2\beta}-\frac{d}{p}}(\mathbb{R}^{d}))\subseteq C_{t,x}^{\alpha-\frac{1}{q},\,\phi_{\gamma-2\beta}-\frac{d}{p}}([0,T]\times\mathbb{R}^{d}).$
Conversely, note that
$\frac{|\mathcal{D}^{L}_{(\tau,0)}u(t,x)|}{|\tau|^{\alpha-\frac{1}{q}}}+\frac{|\mathcal{D}^{L}_{(0,h)}u(t,x)|}{\phi(|h|^{-2})^{-\frac{\gamma}{2}+\beta}|h|^{-\frac{d}{p}}}\leq\sup_{(\tau,h)\neq(0,0)}\frac{\|\mathcal{D}^{L}_{(\tau,h)}u\|_{L_{\infty}(\mathbb{R}^{d}_{T,L,\tau})}}{|\tau|^{\alpha-\frac{1}{q}}+\phi(|h|^{-2})^{-\frac{\gamma}{2}+\beta}|h|^{-\frac{d}{p}}}.$
It remains to prove
$\sup_{t,s\in[0,T]}\frac{|u(t,x)-u(s,x)|}{|t-s|^{\alpha-\frac{1}{q}}}\leq N_{1}\sup_{t\in[0,T]}|u(t,x)|+N_{2}\sup_{t>0}\frac{\omega_{L}(u(\cdot,x),t)}{t^{\alpha-\frac{1}{q}}},$ (3.23)
where $N_{1}=N_{1}(\alpha,L,q,T)$, $N_{2}=N_{2}(\alpha,L,q)$
$\omega_{L}(u(\cdot,x),t):=\sup_{0<|\tau|\leq t}\big{\|}\mathcal{D}^{L}_{(\tau,0)}u(\cdot,x)\big{\|}_{L_{\infty}(\mathrm{T}_{L,\tau})},\qquad\mathrm{T}_{L,\tau}:=\big{\{}s\in[0,T]:\,s+L\tau\in[0,T]\big{\}}.$
Since (3.23) is pointwise in $x$, it suffices to prove it for scalar functions on $[0,T]$. Thus, let $f:[0,T]\to\mathbb{R}$ and set
$\mathcal{D}^{L}_{\tau}f(t):=\sum_{k=0}^{L}(-1)^{L-k}\binom{L}{k}f(t+k\tau),\qquad\omega_{L}(f,t):=\sup_{0<|\tau|\leq t}\big{\|}\mathcal{D}^{L}_{\tau}f\big{\|}_{L_{\infty}(\mathrm{T}_{L,\tau})},$
and
$f]_{C^{\beta}_{L}([0,T])}:=\sup_{t>0}\frac{\omega_{L}(f,t)}{t^{\beta}}.$
Claim 1. Let $a_{j}:=\omega_{L}(f,2^{-j}t)$ and $b_{j}:=\omega_{L+1}\!\left(f,2^{-j}t\right)$. Then,
$a_{0}\leq 2^{L}(1+2^{-LJ})\,a_{J}+L2^{L-1}\sum_{j=0}^{J}(1+2^{-Lj})\,b_{j},\quad\forall\,J\in\mathbb{N}.$ (3.24)
Proof of Claim 1.
Let $I$ be the identity and $T_{\tau}f(t):=f(t+\tau)$. Then
$\mathcal{D}_{\tau}=(I+T_{\tau/2})\mathcal{D}_{\tau/2},$
hence
$\mathcal{D}^{L}_{\tau}-\mathcal{D}^{L}_{\tau/2}$ $=(I+T_{\tau/2})^{L}\mathcal{D}^{L}_{\tau/2}-\mathcal{D}^{L}_{\tau/2}=\sum_{k=1}^{L}\binom{L}{k}(T_{\tau/2}^{k}-2^{-L}I)\,\mathcal{D}^{L}_{\tau/2}$
$=\sum_{k=1}^{L}\binom{L}{k}\sum_{r=0}^{k-1}2^{-\frac{L(k-1-r)}{k}}T_{\tau/2}^{\,r}\,\mathcal{D}^{L+1}_{\tau/2}.$
Therefore,
$\big{\|}\mathcal{D}^{L}_{\tau}f-\mathcal{D}^{L}_{\tau/2}f\big{\|}_{L_{\infty}(\mathrm{T}_{L,\tau})}\leq L2^{L-1}\,\big{\|}\mathcal{D}^{L+1}_{\tau/2}f\big{\|}_{L_{\infty}(\mathrm{T}_{L+1,\tau/2})}.$
Then for integers $0\leq k<J$, define
$S_{k,J}:=\sum_{j=k}^{J-1}2^{-Lj}\Big{(}\mathcal{D}^{L}_{\tau/2^{j}}-\mathcal{D}^{L}_{\tau/2^{j+1}}\Big{)},$
which telescopes to
$S_{k,J}=2^{-Lk}\mathcal{D}^{L}_{\tau/2^{k}}-(2^{L}-1)\sum_{j=k+1}^{J-1}2^{-Lj}\mathcal{D}^{L}_{\tau/2^{j}}-2^{-L(J-1)}\mathcal{D}^{L}_{\tau/2^{J}}.$
Taking sup–norms on the appropriate domains yields
$2^{-Lk}\big{\|}\mathcal{D}^{L}_{\tau/2^{k}}f\big{\|}_{L_{\infty}(\mathrm{T}_{L,\tau/2^{k}})}$ $\leq(2^{L}-1)\sum_{j=k+1}^{J-1}2^{-Lj}\big{\|}\mathcal{D}^{L}_{\tau/2^{j}}f\big{\|}_{L_{\infty}(\mathrm{T}_{L,\tau/2^{j}})}$
$\quad+2^{-L(J-1)}\big{\|}\mathcal{D}^{L}_{\tau/2^{J}}f\big{\|}_{L_{\infty}(\mathrm{T}_{L,\tau/2^{J}})}$
$\quad+L2^{L-1}\sum_{j=k}^{J-1}2^{-Lj}\big{\|}\mathcal{D}^{L+1}_{\tau/2^{j+1}}f\big{\|}_{L_{\infty}(\mathrm{T}_{L+1,\tau/2^{j+1}})}.$
Hence, we obtain the discrete Volterra inequality: for $k=0,1,2,\cdots,J-2$,
$a_{k}\leq 2^{-L(J-1-k)}a_{J}+(2^{L}-1)\sum_{j=k+1}^{J-1}2^{-L(j-k)}a_{j}+L2^{L-1}\sum_{j=k}^{J-1}2^{-L(j-k)}b_{j+1}.$ (3.25)
By summing over $k=0,1,\cdots,J-2$,
$\sum_{k=0}^{J-2}a_{k}$ $\leq a_{J}\sum_{k=0}^{J-2}2^{-L(J-1-k)}+(2^{L}-1)\sum_{k=0}^{J-2}\sum_{j=k+1}^{J-1}2^{-L(j-k)}a_{j}+L2^{L-1}\sum_{k=0}^{J-2}\sum_{j=k}^{J-1}2^{-L(j-k)}b_{j+1}$
$\leq\frac{2^{L}a_{J}}{2^{L}-1}+\sum_{j=1}^{J-1}(1-2^{-Lj})a_{j}+\frac{L2^{L-1}}{2^{L}-1}\sum_{j=0}^{J-1}b_{j+1}.$
Since $a_{k+1}\leq a_{k}$, we have
$\sum_{k=1}^{J-1}a_{k}\leq\sum_{k=0}^{J-2}a_{k}\leq\frac{2^{L}a_{J}}{2^{L}-1}+\sum_{j=1}^{J-1}(1-2^{-Lj})a_{j}+\frac{L2^{L-1}}{2^{L}-1}\sum_{j=0}^{J-1}b_{j+1},$
thus
$\sum_{j=1}^{J-1}2^{-Lj}a_{j}\leq\frac{2^{L}a_{J}}{2^{L}-1}+\frac{L2^{L-1}}{2^{L}-1}\sum_{j=0}^{J-1}b_{j+1}$ (3.26)
Putting $k=0$ in (3.25) and using (3.26) yields (3.24). ∎
Claim 2. For all $t\in(0,T]$,
$\omega_{L}(f,t)\leq N(L)t^{L}\|f\|_{L_{\infty}([0,T])}+N(\alpha,L,q)t^{\alpha-\frac{1}{q}}[f]_{C_{L+1}^{\alpha-\frac{1}{q}}([0,T])}.$ (3.27)
Proof of Claim 2.
Note that for any $J\in\mathbb{N}$,
$1=\frac{1}{\log 2}\int_{2^{-J}t}^{2^{-(J-1)}t}\frac{1}{\lambda}\,\mathrm{d}\lambda,\qquad 2^{-LJ}=\frac{t^{L}}{L}\int_{2^{J}t}^{\infty}\lambda^{-L-1}\,\mathrm{d}\lambda.$
Since $\lambda\mapsto\omega_{L}(f,\lambda)$ is nondecreasing, we have
$(1+2^{-LJ})a_{J}\leq\frac{1}{\log 2}\int_{0}^{2^{-(J-1)}t}\frac{\omega_{L}(f,\lambda)}{\lambda}\mathrm{d}\lambda+\frac{t^{L}}{L}\int_{2^{J}t}^{\infty}\frac{\omega_{L}(f,\lambda)}{\lambda^{L+1}}\mathrm{d}\lambda.$
Using
$\omega_{L}(f,\lambda)\leq\min\left(2^{L}\|f\|_{L_{\infty}([0,T])},\lambda^{\alpha-\frac{1}{q}}[f]_{C_{L}^{\alpha-\frac{1}{q}}([0,T])}\right),$
we obtain, for all $J\in\mathbb{N}$,
$(1+2^{-LJ})a_{J}$
$\leq\frac{[f]_{C_{L}^{\alpha-\frac{1}{q}}([0,T])}}{\log 2}\int_{0}^{2^{-(J-1)}t}\lambda^{\alpha-\frac{1}{q}-1}\mathrm{d}\lambda+\frac{2^{L}t^{L}}{L}\|f\|_{L_{\infty}([0,T])}\int_{2^{J}t}^{\infty}\lambda^{-L-1}\mathrm{d}\lambda$
$\leq N(\alpha,q)t^{\alpha-\frac{1}{q}}2^{-J(\alpha-\frac{1}{q})}[f]_{C_{L}^{\alpha-\frac{1}{q}}([0,T])}+N(L)t^{L}\|f\|_{L_{\infty}([0,T])}\int_{2^{J}t}^{\infty}\lambda^{-L-1}\mathrm{d}\lambda.$
Taking the limit superior as $J\to\infty$ gives
$\limsup_{J\to\infty}(1+2^{-LJ})a_{J}\leq N(L)t^{L}\|f\|_{L_{\infty}([0,T])}\limsup_{J\to\infty}\int_{2^{J}t}^{\infty}\lambda^{-L-1}\mathrm{d}\lambda.$
Since
$\limsup_{J\to\infty}\int_{2^{J}t}^{\infty}\lambda^{-L-1}\mathrm{d}\lambda\leq\int_{1}^{\infty}\lambda^{-L-1}\mathrm{d}\lambda=N(L),$
we have
$\limsup_{J\to\infty}(1+2^{-LJ})a_{J}\leq N(L)t^{L}\|f\|_{L_{\infty}([0,T])}.$ (3.28)
Next, since $\lambda\mapsto\omega_{L+1}(f,\lambda)$ is nondecreasing,
$b_{j+1}=\frac{1}{\log 2}\int_{2^{-(j+1)}t}^{2^{-j}t}\frac{\omega_{L+1}\left(f,2^{-(j+1)}t\right)}{\lambda}\,\mathrm{d}\lambda\leq\frac{1}{\log 2}\int_{2^{-(j+1)}t}^{2^{-j}t}\frac{\omega_{L+1}(f,\lambda)}{\lambda}\,\mathrm{d}\lambda,$
and thus, for any $J\in\mathbb{N}$,
$\sum_{j=0}^{J-1}(1+2^{-Lj})b_{j+1}$ $\leq\frac{1}{\log 2}\sum_{j=0}^{J-1}(1+2^{-Lj})\int_{2^{-(j+1)}t}^{2^{-j}t}\frac{\omega_{L+1}(f,\lambda)}{\lambda}\mathrm{d}\lambda$
$\leq\frac{1}{\log 2}\sum_{j=0}^{J-1}\int_{2^{-(j+1)}t}^{2^{-j}t}\left(1+\left(\frac{2\lambda}{t}\right)^{L}\right)\frac{\omega_{L+1}(f,\lambda)}{\lambda}\mathrm{d}\lambda$
$=\frac{1}{\log 2}\int_{2^{-J}t}^{t}\left(1+\left(\frac{2\lambda}{t}\right)^{L}\right)\frac{\omega_{L+1}(f,\lambda)}{\lambda}\mathrm{d}\lambda$
$\leq\frac{(1+2^{L})}{\log 2}\int_{2^{-J}t}^{t}\frac{\omega_{L+1}(f,\lambda)}{\lambda}\mathrm{d}\lambda.$
Hence,
$\limsup_{J\to\infty}\sum_{j=0}^{J-1}(1+2^{-Lj})b_{j+1}$ $\leq\frac{(1+2^{L})}{\log 2}\int_{0}^{t}\frac{\omega_{L+1}(f,\lambda)}{\lambda}\mathrm{d}\lambda$ (3.29)
$\leq[f]_{C_{L+1}^{\alpha-\frac{1}{q}}([0,T])}\frac{L2^{L-1}}{\log 2}\int_{0}^{t}\lambda^{\alpha-\frac{1}{q}-1}\mathrm{d}\lambda$
$=N(\alpha,L,q)t^{\alpha-\frac{1}{q}}[f]_{C_{L+1}^{\alpha-\frac{1}{q}}([0,T])}.$
Starting from (3.24) and using (3.28) and (3.29), we conclude
$\omega_{L}(f,t)=a_{0}$ $\leq\limsup_{J\to\infty}\left(2^{L}(1+2^{-LJ})a_{J}+L2^{L-1}\sum_{j=0}^{J}(1+2^{-Lj})b_{j}\right)$
$\leq N(L)t^{L}\|f\|_{L_{\infty}([0,T])}+N(\alpha,L,q)t^{\alpha-\frac{1}{q}}[f]_{C^{\alpha-\frac{1}{q}}_{L+1}([0,T])},$
which is (3.27). ∎
Now, we prove (3.23). Dividing both sides of (3.27) by $t^{\alpha-\frac{1}{q}}$ and taking the supremum over $t\in(0,T]$, we obtain
$[f]_{C^{\alpha-\frac{1}{q}}_{L}([0,T])}\leq N(L)\,T^{L-\alpha+\frac{1}{q}}\|f\|_{L_{\infty}([0,T])}+N(\alpha,L,q)\,[f]_{C^{\alpha-\frac{1}{q}}_{L+1}([0,T])}.$ (3.30)
Since
$\sup_{t,s\in[0,T]}\frac{|f(t)-f(s)|}{|t-s|^{\alpha-\frac{1}{q}}}\leq\sup_{t>0}\frac{\omega_{1}(f,t)}{t^{\alpha-\frac{1}{q}}}=:[f]_{C^{\alpha-\frac{1}{q}}_{1}([0,T])},$
(3.30) yields
$\sup_{t,s\in[0,T]}\frac{|f(t)-f(s)|}{|t-s|^{\alpha-\frac{1}{q}}}$ $\leq[f]_{C^{\alpha-\frac{1}{q}}_{1}([0,T])}$
$\leq N(L)[f]_{C^{\alpha-\frac{1}{q}}_{2}([0,T])}+N(\alpha,L,q)\|f\|_{L_{\infty}([0,T])}$
$\leq\cdots\leq N_{1}\|f\|_{L_{\infty}([0,T])}+N_{2}[f]_{C^{\alpha-\frac{1}{q}}_{L}([0,T])}.$
Finally, applying this with $f(\cdot)=u(\cdot,x)$ gives (3.23). This completes the proof. ∎
Remark 3.10.
The choice of $L$ in Definition 3.8 is immaterial for the definition of
$C_{t,x}^{\alpha-\frac{1}{q},\,\phi_{\gamma-2\beta}-\frac{d}{p}}([0,T]\times\mathbb{R}^{d}).$
Indeed, by Proposition 3.9, the parameter $L$ appears only in the definition of $C^{\phi_{\gamma-2\beta}-\frac{d}{p}}(\mathbb{R}^{d})$. Moreover, by [11, Theorem 1.5],
$|u|_{C^{\phi_{\gamma-2\beta}-\frac{d}{p}}(\mathbb{R}^{d})}\simeq\|u\|_{B_{\infty,\infty}^{\phi_{\gamma,\beta,p,d}}(\mathbb{R}^{d})},$
where $\phi_{\gamma,\beta,p,d}(\lambda):=\phi(\lambda^{2})^{\frac{\gamma}{2}-\beta}\lambda^{-\frac{d}{p}}$. It is straightforward to verify that the definition of
$B_{\infty,\infty}^{\phi_{\gamma,\beta,p,d}}(\mathbb{R}^{d})$
does not depend on the choice of $L$.
@@
Let $\delta_{0}\in(0,1]$, $\gamma\in(0,\infty)$, $p\in[2,\infty)$, $q\in(2,\infty)$, and suppose Assumption 1.12 ($\delta_{0}$) holds. If
$\frac{1}{q}\leq\alpha<\beta\leq\frac{1}{2},\qquad\delta_{0}(\gamma-2\beta)-\frac{d}{p}>0,$
then the space $\mathcal{H}_{p,q}^{\phi,\gamma}(T)$ is continuously embedded into $L_{q}\left(\Omega;\,C_{t,x}^{\alpha-\frac{1}{q},\,\phi_{\gamma-2\beta}-\frac{d}{p}}([0,T]\times\mathbb{R}^{d})\right)$. Equivalently, there exists a constant $N=N(d,N_{0},N_{1},N_{2},p,q,T)$ such that for all $u\in\mathcal{H}_{p,q}^{\phi,\gamma}(T)$,